# Taming Latent Diffusion Models to See in the Dark

### 10 Dec 2023 v2

@@ 4.1 Low-frequency Content Generation
Figure 4: The low-frequency sub-band, extracted using 2D discrete wavelet transforms (DWT), reveals clearer structural information compared to the low-light input. On the other hand, merely rescaling the input to align with the latent size of the LDM indicates that the information remains predominantly influenced by the noise. The result “Ours (Rescaling)” referring to merely rescaling the input for taming the LDM shows over-smooth and poorly-structured results due to the negative impact of the noise.
Employing a series of DWT not only extracts the low-frequency sub-band (LL) but also allows us to effectively align the input information with the latent size of the LDM. Compared to directly rescaling the input RAW image, feeding the low-frequency sub-band (LL) to the LDM enables the model to concentrate on generating high-correlation structural content while reducing the negative impact of the noise, as shown in Fig. 4.
By initializing a latent representation $Z_{T}\sim\mathcal{N}(0,I)$, we aim at steering the denoising process in the LDM by inserting a taming module at each layer. After denoising $T$ timesteps, we obtain a clean latent representation $Z_{0}$ showing a high correlation to the low-frequency sub-band (LL).
@@
For the taming modules of the LDM, we employ a similar time-aware guidance strategy as discussed in [27]. However, instead of deriving the structural information by encoding the input image with the encoder $\mathcal{E}$ as in  [27], we directly utilize the low-frequency sub-band (LL) to provide our structural information. This liberates us from the substantial computational cost of encoding.
We insert a taming module at each layer in the LDM. The taming module consists of a set of convolutional layers and an SFT layer for mapping the low-frequency sub-band (LL) to a pair of affine transformation parameters:
@@
where $l$ denotes the layer number in the LDM and t denotes the timestep. Then, the feature $\textbf{\text{F}}^{\text{LDM}}_{l}$ at the layer $l$ in the LDM is modulated by
$\hat{\textbf{\text{F}}}^{\text{LDM}}_{l}=(1+\alpha_{l})\odot\textbf{\text{F}}^{\text{LDM}}_{l}+\beta_{l}.$ (2)
After taming the LDM to generate the latent representation $Z_{0}$, an sRGB can be directly obtained by feeding it to the decoder $\mathcal{D}$. As shown in Fig. 7, such sRGB image presents well-exposed and high-correlation structural content but inaccurate color and over-generated details. This inspires us to tame the decoder $\mathcal{D}$ during upsampling to guarantee exact illumination and precise details in the final sRGB image.
@@ 4.2 High-frequency Detail Maintenance
This high-frequency “taming” aims to leverage the noise-free and fine-grained generative prior in the decoder to produce final sRGB images with enhanced details and devoided noise.
Similar to the taming process in the LDM, we insert a taming module at each upsampling layer in the decoder $\mathcal{D}$. For taming the decoder $\mathcal{D}$, not only the high-frequency subbands are needed but also the low-frequency. There are two key reasons for that:
@@
4.3 Training Objective
For training the taming modules of the LDM, the ground-truth sRGB image $Y$ is first encoded into a latent representation $Z_{0}$ via the encoder $\mathcal{E}$. Then, the latent representation is corrupted with noise using the standard forward diffusion schedule [10]: $Z_{0}\overset{\epsilon}{\rightarrow}Z_{t}$, where the noise $\epsilon\sim\mathcal{N}(0,I)$ and the timestep $t\in\left[1,T\right]$. Since the LDM serves to predict the noise from the noised representation, the training objective is
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{Z_{0},\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}(Z_{t},t,\tau^{\text{LDM}}_{\phi}(\text{LL}))\right\|^{2}_{2}\right],$ (3)
where $\epsilon_{\theta}(\cdot)$ denotes the LDM and $\tau^{\text{LDM}}_{\phi}(\cdot)$ denotes taming modules.
For training the taming modules of the decoder $\mathcal{D}$, a clean representation $\hat{Z_{0}}$ is generated by our tamed LDM for each low-light input image and the training objective is
@@ 4.3 Training Objective
Note that all parameters from the LDM and decoder $\mathcal{D}$ are frozen and only our proposed taming modules are trainable.
Figure 5: The qualitative evaluations on SID-Sony [4] (1-4 rows) and ELD-Sony [32] (5-6 rows) datasets (Zoom-in for best view). The input images are amplified and post-processed with the same operations as ELD [32]. The proposed method shows notable superiority in recovering structural information and enhancing details in extremely dark and noisy regions. In the $\text{5}^{\text{th}}$ row, the proposed method recovers the text much more effectively than previous works.
Table 1: The quantitative comparison in extremely challenging scenarios from different datasets. The best score is denoted as boldfaced. The proposed method significantly outperforms previous works according to perceptual metrics, including LPIPS [40], NIMA [24] and CLIP-IQA [26]. It indicates the sRGB results obtained from the proposed method exhibit significantly clearer global structure and more natural details compared to those from previous works.
@@ Methods
PSNR $\uparrow$
@@
NIMA $\uparrow$
CLIP-IQA $\uparrow$
SID-Sony [4] $\times$300 SID [4] 27.1862 0.8219 0.3906 3.0481 0.2279
ELD [32] 27.4026 0.8271 0.3876 3.0346 0.2342
LED [14] 26.3541 0.8228 0.4168 3.1831 0.2688
ExposureDiff [30] 28.0544 0.8383 0.3892 2.9931 0.2519
Ours 27.2857 0.8113 0.2749 3.5583 0.3248
ELD-Sony [32] $\times$200 SID [4] 28.9735 0.7990 0.2618 2.9780 0.2057
ELD [32] 29.0467 0.8063 0.2585 3.1798 0.2504
LED [14] 30.4692 0.8769 0.2058 3.1718 0.2883
ExposureDiff [30] 30.7530 0.8582 0.1975 3.0221 0.2915
Ours 29.8284 0.8906 0.2035 3.4397 0.3448
LRD [38] -3EV LRD [38] 29.4312 0.8784 0.2198 3.1985 0.3078
Ours 29.0672 0.8889 0.1858 3.6180 0.3862
DRV [5] [$\times$100, $\times$300] Retinexformer [2] 29.3872 0.9030 0.1460 4.7620 0.3326
Ours 28.4360 0.8900 0.1304 4.7839 0.4318
Figure 6: The qualitative evaluations on LRD [38] (the first row) and DRV [5] (the second row) datasets (Zoom-in for best view).
5 Experiments
5.1 Implementation Details
The proposed method is evaluated on four widely-used datasets for the LLIE task, including SID-Sony [4], ELD-Sony [32], LRD [38], and DRV [5]. The train and test subsets follow the same strategy as previous works. Since the proposed method is fed with a RAW input and outputs an sRGB image, we convert the RAW output of previous works with the same ISP pipeline (e.g., rawpy [^1] https://pypi.org/project/rawpy/) to guarantee fair comparisons.
During training, the taming modules of the LDM are trained for 100 epochs while the taming modules of the decoder are trained for 30 epochs. By initially taming the LDM, it gains the capability to generate the clean latent representation $Z_{0}$ for each low-light training image. This generated latent representation is subsequently utilized to train the decoder taming modules. The training process is performed on 512 $\times$ 512 resolution with 4 NVIDIA GeForce RTX 3090 GPUs. During inference, the LDM follows the DDIM sampling strategy [23] with 200 timesteps.
5.2 Quantitative Evaluation
In this work, we mainly focus on the perceptual quality of results which are enhanced from extremely dark and noisy images. Besides distortion metrics (PSNR and SSIM), we also adopt LPIPS [40], NIMA [24], and CLIP-IQA [26] for our quantitative evaluations. LPIPS [40] and NIMA [24] are two widely-used reference-based and non-reference-based neural perceptual metrics. CLIP-IQA [26] is a state-of-the-art non-reference-based image quality perception assessment method. It is based on a large model, CLIP [19], to assess an image considering quality and abstract perceptions.
According to Table  1, the proposed method outperforms other methods in terms of non-reference-based perceptual metrics, including NIMA [24] and CLIP-IQA [26]. It indicates the sRGB results obtained from the proposed method exhibit significantly more natural details compared to those from previous works. The proposed method does not achieve the best but still competitive performance in terms of PSNR metric. There are two potential reasons: 1) Some misalignment exists between the low-light input and long-exposure ground truth; 2) Noticeable noise still remains in the long-exposure ground truth [38]. The structural quality assessment based on SSIM and LPIPS [40] indicates the proposed method has substantial advantages in enhancing low-frequency information, especially structural content.
Figure 7: The visualization of the ablation study. Different taming strategies (rescaling and DWT) and taming targets (LDM and decoder $\mathcal{D}$) are combined in this study. Taming the LDM by rescaling the input information tends to result in an ambiguous global structure. In contrast, the taming result using 2D discrete wavelet transforms (DWT) exhibits distinct structural content. Moreover, taming the decoder $\mathcal{D}$ results in improved color distribution and the enhancement of details.
5.3 Qualitative Evaluation
Fig. 5 and Fig. 6 show the visual comparison with other state-of-the-art methods on the SID-Sony [4], ELD-Sony [32], LRD [38] and DRV [5] datasets. Images captured in extremely low-light conditions require significant amplification ratios ($\times$300 or -3 EV). These amplified images are plagued by acute noise and color distortion, making most details challenging to discern with the human eye. Compared with previous works, the proposed method shows notable superiority in recovering structural information and enhancing details in noisy regions. In the $\text{2}^{\text{nd}}$ row of Fig. 5, the results produced by previous works tend to be over-smooth due to the excessive denoising while the proposed method present sharp and clean leaves without color bias. In the $\text{5}^{\text{nd}}$ row of Fig. 5, previous works struggle to produce meaningful text information when portions of it are obscured by intense noise in the low-light image. The proposed method, which leverages the potent generative prior from pre-trained Stable Diffusion, exhibits a significantly improved reconstruction of the text.
5.4 Ablation Study
Table 2: The ablation study of the proposed method on ELD-Sony dataset [32]. Different taming strategies (rescaling and DWT) and taming targets (LDM and decoder $\mathcal{D}$) are combined in this study. The best score is denoted as boldfaced.
LDM
Decoder $\mathcal{D}$
PSNR $\uparrow$
SSIM $\uparrow$
LPIPS $\downarrow$
Rescaling - 26.3223 0.7544 0.5423
27.3342 0.8166 0.2934
DWT - 26.9322 0.7652 0.4904
29.8284 0.8906 0.2035
@@ 5.4 Ablation Study
As we mentioned before, 2D discrete wavelet transforms (DWT) not only extract sub-bands for the taming but also downsample the input information to align with the latent size in LDM. Instead of DWT, there are two other intuitive ways for downsampling: rescaling and CNN-based downsampling. In this study, we forego discussion on CNN-based downsampling due to its substantial computational expense. As shown in Fig. 4, simply rescaling the input (e.g., using Bicubic Interpolation) results in the loss of many details and causes the small-scale input to be overwhelmed by noise. Taming the LDM with such input could result in an ambiguous global structure, potentially leading to low-quality sRGB images after decoding. A qualitative comparison between rescaling and DWT taming strategies is shown in Fig. 7.
@@
The taming modules of the LDM hold significant importance as they steer the generating process toward generating specified latent representations. There is no need to conduct an ablation study on whether taming the LDM since the LDM without taming generates unexpected latent representation which is not related to the low-light input.
The fact that motivates us to tame the decoder $\mathcal{D}$ is that although the low-frequency sub-band (LL) helps the proposed method to tame the LDM for generating a high-correlation latent representation, the sRGB image decoded from this representation by a decoder lacking taming modules, shows much over-generated and irrelevant details.
@@ 6 Conclusion
To improve the LLIE performance in extremely dark and noisy images, this work extricates itself from collecting real-world data or carefully designing noise models for synthesizing challenging data. Instead, we propose to tame a pre-trained diffusion model and explore its generative priors for achieving the LLIE task. Our designed taming modules are fed with the low-light input information, and modulate intermediate features within the diffusion model, which allows us to steer the generating process. With the observation of dedicated generative priors in different portions of the diffusion model, we propose to utilize 2D discrete wavelet transforms to divide the LLIE task into two parts: the low-frequency content generation and the high-frequency detail maintenance. Benefiting from the abundant and dedicated generative priors in the diffusion model, extensive experiments demonstrate that taming the diffusion model allows us to achieve the optimized balance between image fidelity and perceptual quality, leading to state-of-the-art LLIE performance.

### 20 Mar 2024 v3

@@ 3.2 Low-frequency Content Generation
Figure 4: The low-frequency sub-band, extracted using 2D discrete wavelet transforms (DWT), reveals clearer structural information compared to the low-light input. On the other hand, merely resizing the input to align with the latent size of the UNet indicates that the information remains predominantly influenced by the noise. The result “Ours (Resize)” referring to merely resizing the input for taming the latent diffusion model shows over-smooth and poorly-structured results due to the negative impact of the noise.
Employing a series of DWT not only extracts the low-frequency sub-band (LL) but also allows us to efficiently align the input information with the latent size of the UNet. Compared to directly resizing the input RAW image, feeding the low-frequency sub-band (LL) to the UNet enables the model to concentrate on generating high-correlation structural content while reducing the negative impact of the noise, as shown in Fig. 4.
By initializing a latent representation $Z_{T}\sim\mathcal{N}(0,I)$, we aim at steering the denoising process in the UNet by inserting a taming module at each layer. After denoising $T$ timesteps, we obtain a clean latent representation $Z_{0}$ showing a high correlation to the low-frequency sub-band (LL).
@@
For the taming modules of the UNet, we employ a similar time-aware guidance strategy as discussed in [26]. However, instead of deriving the structural information by encoding the input image with the encoder $\mathcal{E}$ as in [26], we directly utilize the low-frequency sub-band (LL) to provide our structural information. This liberates us from the substantial computational cost of encoding, as shown in Table 1.
Table 1: A computational comparison between the encoder $\mathcal{E}$ adopted in [26] and our adopted DWT.
Input size
Time (s)
Params (M)
MACs (M)
Finetune for RAW
Encoder $\mathcal{E}$ $512\times 512$ 2.09 34.16 566,371
DWT $512\times 512$ 0.02 0 -
We insert a taming module at each layer in the UNet. The taming module consists of a set of convolutional layers and an SFT layer for mapping the low-frequency sub-band (LL) to a pair of affine transformation parameters:
@@
where $l$ denotes the layer number in the UNet and t denotes the timestep. Then, the feature $\textbf{\text{F}}^{\text{UNet}}_{l}$ at the layer $l$ in the UNet is modulated by
$\hat{\textbf{\text{F}}}^{\text{UNet}}_{l}=(1+\alpha_{l})\odot\textbf{\text{F}}^{\text{UNet}}_{l}+\beta_{l}.$ (2)
After taming the UNet to generate the latent representation $Z_{0}$, an sRGB can be directly obtained by feeding it to the decoder $\mathcal{D}$. As shown in Fig. 7, such sRGB image presents well-exposed and high-correlation structural content but inaccurate color and over-generated details. This inspires us to tame the decoder $\mathcal{D}$ during upsampling to guarantee exact illumination and precise details in the final sRGB image.
@@ 3.3 High-frequency Detail Maintenance
This high-frequency “taming" aims to leverage the noise-free and fine-grained generative prior in the decoder to produce final sRGB images with enhanced details and devoided noise.
Similar to the taming process in the UNet, we insert a taming module at each upsampling layer in the decoder $\mathcal{D}$. For taming the decoder $\mathcal{D}$, not only the high-frequency subbands are needed but also the low-frequency. There are two key reasons for that:
@@
3.4 Training Objectives
For training the taming modules of the UNet, the ground-truth sRGB image $Y$ is first encoded into a latent representation $Z_{0}$ via the encoder $\mathcal{E}$. Then, the latent representation is corrupted with noise using the standard forward diffusion schedule [7]: $Z_{0}\overset{\epsilon}{\rightarrow}Z_{t}$, where the noise $\epsilon\sim\mathcal{N}(0,I)$ and the timestep $t\in\left[1,T\right]$. Since the UNet serves to predict the noise from the noised representation, the training objective is
$\mathcal{L}_{\text{UNet}}=\mathbb{E}_{Z_{0},\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}(Z_{t},t,\tau^{\text{UNet}}_{\phi}(\text{LL}))\right\|^{2}_{2}\right],$ (3)
where $\epsilon_{\theta}(\cdot)$ denotes the UNet and $\tau^{\text{UNet}}_{\phi}(\cdot)$ denotes taming modules.
For training the taming modules of the decoder $\mathcal{D}$, a clean representation $\hat{Z_{0}}$ is generated by our tamed UNet for each low-light input image and the training objective is
@@ 3.4 Training Objectives
Note that all parameters from the UNet and decoder $\mathcal{D}$ are frozen and only our proposed taming modules are trainable.
Figure 5: The qualitative evaluations on SID-Sony [2] and ELD-Sony [30] datasets (Zoom-in for best view). The proposed method shows notable superiority in recovering structural information and enhancing details in extremely dark and noisy regions. ($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
Table 2: The quantitative comparison in extremely challenging scenarios from different datasets. The proposed method achieves state-of-the-art performance according to perceptual metrics, including LPIPS [39] and NIMA [25]. It indicates the sRGB results obtained from the proposed method exhibit significantly clearer global structure and more natural details compared to those from previous works. The best score is denoted as boldfaced.($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
@@ Methods
LPIPS $\downarrow$
NIMA $\uparrow$
SSIM $\uparrow$
SID-Sony [2] $\times$300 ${}^{\dagger}$ [2] 0.3906 3.0481 0.8219
${}^{\ast}$ [30] 0.3876 3.0346 0.8271
${}^{\ast}$ [12] 0.4168 3.1831 0.8228
${}^{\dagger}$ [4] 0.4794 2.8077 0.7749
${}^{\dagger}$ [11] 0.4176 2.7646 0.8448
${}^{\dagger}$ 0.2749 3.5583 0.8113
ELD-Sony [30] $\times$200 ${}^{\dagger}$ [2] 0.2618 2.9780 0.7990
${}^{\ast}$ [30] 0.2585 3.1798 0.8063
${}^{\ast}$ [12] 0.2058 3.1718 0.8769
${}^{\dagger}$ [4] 0.3028 2.8887 0.7708
${}^{\dagger}$ [11] 0.2172 2.9123 0.8605
${}^{\dagger}$ 0.2035 3.4397 0.8906
LRD [38] -3EV ${}^{\ast}$ [38] 0.2198 3.1985 0.8784
${}^{\dagger}$ 0.1858 3.6180 0.8889
Figure 6: The qualitative evaluations on LRD [38] dataset (Zoom-in for best view).
4 Experiments
4.1 Implementation Details
The proposed method is evaluated on three widely-used datasets for the LLIE task, including SID-Sony [2], ELD-Sony [30], and LRD [38]. The train and test subsets follow the same strategy as previous works. Given the ground-truth RAW images from these datasets, we obtain their corresponding sRGB images through the open-source ISP pipeline, rawpy***https://pypi.org/project/rawpy/. The same ISP pipeline is also adopted for the outputs of our RAW-to-RAW baselines.
During training, the taming modules of the UNet are trained for 100 epochs while the taming modules of the decoder are trained for 30 epochs. By initially taming the UNet, it gains the capability to generate the clean latent representation $Z_{0}$ for each low-light training image. This generated latent representation is subsequently utilized to train the decoder taming modules. The training process is performed on 512 $\times$ 512 resolution with 4 NVIDIA GeForce RTX 3090 GPUs. During inference, the UNet follows the DDIM sampling strategy [24] with 200 timesteps.
4.2 Baselines
To demonstrate the superiority of the proposed LDM-ISP on enhancing ISP for low-light images, we first conduct a comparison with strong RAW-to-sRGB LLIE methods. To further highlight our denoising performance, we also compare the proposed method with RAW-to-RAW methods. The state-of-the-art methods compared with the proposed LDM-ISP are:
  • RAW-to-sRGB methods: SID [2], MCR [4] and DNF [11].
  • RAW-to-RAW methods: ELD [30], LED [12] and LRD [38].
4.3 Quantitative Evaluation
In this work, we mainly focus on the perceptual quality of sRGB results which are enhanced from extremely dark and noisy RAW images. Besides the metric SSIM, we also adopt LPIPS [39] and NIMA [25]for our quantitative evaluations. LPIPS [39] and NIMA [25] are two widely-used reference-based and non-reference-based neural perceptual metrics.
According to Table  2, the proposed method outperforms other methods in terms of neural perceptual metrics, LPIPS [39] and NIMA [25]. To be specific, the proposed method achieves 0.1127 and 0.0023 improvements in LPIPS [39] than the previous sate-of-the-art methods on SID-Sony [2] and ELD-Sony [30] subsets, respectively. Meanwhile, there are 0.3752 and 0.2599 improvements in NIMA [25] on two subsets, respectively. It indicates the sRGB results obtained from the proposed method exhibit significantly more natural details compared to those from previous works. The structural quality assessment based on SSIM and LPIPS [39] indicates the proposed method has substantial advantages in enhancing low-frequency information, especially structural content. Note that the proposed method not only achieves state-of-the-art ISP performance compared to RAW-to-sRGB methods but also presents remarkably better denoising performance than RAW-to-RAW methods, which is a significant challenge to previous RAW-to-sRGB methods.
Figure 7: The visualization of the ablation study. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. Taming the UNet by resizing the input information tends to result in an ambiguous global structure. In contrast, the taming result using 2D discrete wavelet transforms (DWT) exhibits distinct structural content. Moreover, taming the decoder $\mathcal{D}$ results in improved color distribution and the enhancement of details.
4.4 Qualitative Evaluation
Fig. 5 and Fig. 6 show the visual comparison with other state-of-the-art methods on the SID-Sony [2], ELD-Sony [30], and LRD [38] datasets. Images captured in extremely low-light conditions require high amplification ratios ($\times$300 or -3 EV). These amplified images are plagued by acute noise and color distortion, making most details challenging to discern with the human eye. Compared with previous works, the proposed method shows notable superiority in recovering structural information and enhancing details in noisy regions. In the $\text{2}^{\text{nd}}$ column of Fig. 5, the results produced by previous works tend to be over-smooth due to the excessive denoising while the proposed method present sharp and clean leaves without color bias. In the $\text{5}^{\text{nd}}$ column of Fig. 5, previous works struggle to produce meaningful text information when portions of it are obscured by intense noise in the low-light image. The proposed method, which leverages the potent generative prior from pre-trained Stable Diffusion, exhibits a significantly improved reconstruction of the text.
4.5 Ablation Study
Table 3: The ablation study of the proposed method on ELD-Sony dataset [30]. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. The best score is denoted as boldfaced.
UNet
Decoder $\mathcal{D}$
@@ 4.5 Ablation Study
Resize - 0.7544 0.5423
0.8233 0.2834
DWT - 0.7652 0.4904
0.8906 0.2035
@@
As we mentioned before, 2D discrete wavelet transforms (DWT) not only extract sub-bands for the taming but also downsample the input information to align with the latent size in the UNet. Instead of DWT, there are two other intuitive ways for downsampling: rescaling and CNN-based downsampling. In this study, we forego discussion on CNN-based downsampling due to its substantial computational expense. As shown in Fig. 4, simply resizing the input (e.g., using Bicubic Interpolation) results in the loss of many details and causes the small-scale input to be overwhelmed by noise. Taming the UNet with such input could result in an ambiguous global structure, potentially leading to low-quality sRGB images after decoding. A qualitative comparison between resizing and DWT taming strategies is shown in Fig. 7.
@@
The taming modules of the UNet hold significant importance as they steer the generating process toward generating specified latent representations. There is no need to conduct an ablation study on whether taming the UNet since the UNet without taming generates unexpected latent representation which is not related to the low-light input.
The fact that motivates us to tame the decoder $\mathcal{D}$ is that although the low-frequency sub-band (LL) helps the proposed method to tame the UNet for generating a high-correlation latent representation, the sRGB image decoded from this representation by a decoder lacking taming modules, shows much over-generated and irrelevant details.
@@ 5 Conclusion
To enhance neural ISP for the LLIE performance in extremely dark and noisy images, this work extricates itself from collecting more real-world data for training learning-based methods and designing more sophisticated network architecture. Instead, we propose to tame a pre-trained diffusion model and explore its generative priors for achieving the LLIE task. Our designed taming modules are fed with the low-light input information, and modulate intermediate features within the diffusion model, which allows us to steer the generating process. With the observation of dedicated generative priors in different portions of the diffusion model, we propose to utilize 2D discrete wavelet transforms to divide the LLIE task into two parts: the low-frequency content generation and the high-frequency detail maintenance. Benefiting from the abundant and dedicated generative priors in the diffusion model, extensive experiments demonstrate that taming the diffusion model allows us to achieve the optimized balance between image fidelity and perceptual quality, leading to state-of-the-art LLIE performance.